OpenStudy (anonymous):

I have a new question to be answered :D FInd the vertex of the graph of the following equation: y= x^2+8x-12. A) (4,-28) B) (-4,-28) C) (-4,4) D) (4,4)

OpenStudy (rational):

change the given equation into vertex form by completing the square

OpenStudy (triciaal):

vertex form a(x-h)^2 + k from general form ax^2 + bx + c= 0 h = -b/2a the x-coordinate of the vertex k = f(x)

OpenStudy (anonymous):

its either b or c

OpenStudy (anonymous):

is it c?

OpenStudy (anonymous):

When I graphed it on my graphing calculator I got -4, -28 as the vertex.

OpenStudy (anonymous):

thank you so B is the correct answer xP